The edge of a cube was found to be 1.50 plus or minus 0.05 what is its surface area and its volume.

1
Expert's answer
2022-03-01T17:33:25-0500

1. The area is "A = 6\cdot 1.50^2 = 13.5". The error is (see https://www.geol.lsu.edu/jlorenzo/geophysics/uncertainties/Uncertaintiespart2.html):


"\Delta A = A\cdot 6\cdot 2\cdot \dfrac{0.05}{1.5}\approx 5"

Thus


"A = 14 \pm5"

2. The volume is "V = 1.50^3 = 3.375". The error is:


"\Delta V = V\cdot 3\cdot \dfrac{0.05}{1.5} \approx 0.3"

Thus, the volume is:


"V = 3.4\pm 0.3"